# The SNO+ Collaboration

### v2

@@ I Introduction
Cosmogenic neutron production arises from a variety of processes. The muons themselves can produce neutrons through direct interactions with nuclei, known as spallation. As muons propagate through matter, they lose energy via ionisation and radiative processes, producing secondary particles that induce electromagnetic and hadronic showers. Neutrons are generated both in photonuclear interactions within the electromagnetic showers, and through spallation processes involving secondary hadrons such as protons, pions, and other nuclear fragments [wang_predicting_2001].
The number of neutrons associated with a muon depends on the muon’s energy. Many underground experiments, with varying average muon energies, have measured the cosmogenic neutron yield. The relation between muon energy and neutron production was most recently demonstrated by Daya Bay [an_cosmogenic_2018], using liquid scintillator-based measurements [hertenberger_muon-induced_1995, boehm_neutron_2000, blyth_measurement_2016, abe_production_2010, collaboration_measurement_1999, bellini_cosmogenic_2013].
The average muon energy in an underground laboratory is dictated by the overburden. To date, the neutron yield measurement with the highest average muon energy was made by the Liquid Scintillator Detector (LSD) experiment [aglietta_neutron_1989]. Measurements using muons of comparable average energies have since been made by the Sudbury Neutrino Observatory (SNO) experiment [aharmim_cosmogenic_2019] and at the China Jinping Underground Laboratory [zhao_measurement_2022]. The large overburden at SNO+ means the measurement in this paper will contribute to the high energy muon data.
The medium through which a muon propagates could also play a role in the number of neutrons produced. Many measurements of the neutron yield have been performed with liquid scintillator [an_cosmogenic_2018, aglietta_neutron_1989, zhao_measurement_2022, abe_production_2010, boehm_neutron_2000, hertenberger_muon-induced_1995, blyth_measurement_2016, bellini_cosmogenic_2013, collaboration_measurement_1999]. SNO provided results using both pure heavy water ($\textup{D}_{2}\textup{O}$) and NaCl-loaded heavy water. Both these results were consistent with each other. The first measurement using $\textup{H}_{2}\textup{O}$ came from Super-Kamiokande (SK), which used gadolinium-loaded water. The measurements by SK [collaboration_measurement_2023], in water, and KamLAND [abe_production_2010], in scintillator, were made using the same muon flux, providing a unique comparison of neutron production in different media. Their results were found to be consistent with each other when normalised by medium density [collaboration_measurement_2023].
@@ II Experimental Set-up
SNO+ is a neutrino experiment that will search for neutrinoless double beta decay. Other physics goals of SNO+ include measurements of reactor, geo- and solar neutrinos. The detector infrastructure was inherited from the SNO experiment. While the SNO+ detector is currently filled with liquid scintillator, it was initially filled with ultra-pure water. During this period, known as the water phase, the electronics from SNO were re-commissioned and the background rates evaluated.
The water phase was used as an opportunity to run the detector as a low-background water Cherenkov experiment. The water phase lasted from May 2017 to July 2019. Analyses from this phase include measurements of the $^{8}\textup{B}$ solar neutrino flux [collaboration_measurement_2019, collaboration_measurement_2024], nucleon decay searches [collaboration_search_2019, collaboration_improved_2022], and the first evidence of detecting reactor antineutrinos in a water Cherenkov detector [collaboration_evidence_2023].
@@ II.1 The SNO+ Detector
The SNO+ detector, shown in Figure 1, is located $2\,\textup{km}$ ($6010\,\textup{m.w.e.}$) underground at SNOLAB [lawson_snolab_2013, smith_snolab_2012]. An acrylic vessel (AV), of radius $6\,\textup{m}$ and thickness $5.5\,\textup{cm}$, contains the target medium that defines the phase. During the water phase, the AV was filled with $905\,\textup{tonnes}$ of ultra-pure water (UPW). A $7\,\textup{m}$ tall cylindrical neck is connected to the top of the AV, through which the AV can be filled and calibration sources can be deployed. At the top and facing down the neck are $4$ neck photomultiplier tubes (PMTs).
@@
Figure 1: Diagram of the SNO+ detector with labelled components [collaboration_sno_2021].
Part way through the water phase in September 2018, a cover gas system was installed to reduce the rate of $^{222}\textup{Rn}$ induced backgrounds in the AV. This system uses nitrogen gas, which acts as a barrier to radon in the air above the liquid in the neck. A detailed description of the SNO+ detector and its cover gas systems can be found in [collaboration_sno_2021].
@@ II.2 Event Triggers
Each PMT has an adjustable channel threshold. When this threshold is exceeded, a hit is registered for that PMT, and a fixed-amplitude pulse of width $89\,\textup{ns}$ is produced. These pulses are summed across all PMTs. An event is recorded if the sum of the PMT pulses surpasses the event trigger threshold. For the water phase, the trigger threshold was set to $7$ pulses, approximately equivalent to a $1.4\,\textup{MeV}$ electron at the centre of the detector [collaboration_measurement_2020]. The trigger system is modelled in SNO+ simulations, as described in Section III. The resulting data is handled by the data acquisition systems, which prepare it for offline analysis.
The trigger efficiency is the probability that an event will trigger the detector, given a number of PMT hits. Using simulation, it was found that the trigger efficiency is $100\,\%$ at and above $8$ PMT hits. Below $8$ PMT hits, the trigger efficiency decreases to zero. As a consequence, the corresponding total efficiency for triggering on a $2.2\,\textup{MeV}$ gamma ray, the signal for neutron capture, is approximately $50\,\%$ [collaboration_measurement_2020].
The distribution of the summed PMT pulses can depend on the nature of the event. If the PMTs involved in an event all register hits at similar times, the sum of the pulses will form a sharper, more narrow peak than if the PMT hits were more spread over time. Close to the threshold, this event-dependence can affect whether an event triggers the detector or not. To minimise the impact of the trigger’s dependency on the PMT hit distribution, the trigger efficiency is calculated as a function of ‘in-time hits’, the number of PMT hits that arrive within an $89\,\textup{ns}$ time window, centred on the peak of the summed PMT pulses.
@@ II.3 Am-Be Calibration
The results of this calibration were presented in [collaboration_measurement_2020]. The efficiency for detecting neutrons was found to be approximately $50\,\%$ near the centre of the detector. The neutron capture time was also measured to be $202.53^{+0.87}_{-0.76}\,\mu\textup{s}$. This neutron detection efficiency is relatively high for a water Cherenkov detector, allowing for the unique measurement of cosmogenic neutrons in UPW.
@@ III Simulation
For this analysis, Monte Carlo (MC) simulations were performed for characterising the reconstruction of muons and neutrons, and measuring the efficiency for detecting and selecting neutrons. Simulations in SNO+ are handled by RAT, a software package used for offline analysis [noauthor_rat_nodate]. All low level physics simulation and particle tracking is performed using the GEANT4 (version 10.00.p04) software package [agostinelli_geant4simulation_2003], which has been incorporated into RAT, along with the detector geometry and trigger performance. Calibrations have been performed in order to tune and verify the full SNO+ simulation model.
@@
where $\theta$ is the angle from zenith from which the muon arrives. The parameters used are $I_{1}=(8.60\pm 0.53)\times 10^{-6}\,\textup{sec}^{-1}\textup{cm}^{-2}\textup{sr}^{-1}$, $I_{2}=(0.44\pm 0.06)\times 10^{-6}\,\textup{sec}^{-1}\textup{cm}^{-2}\textup{sr}^{-1}$, $\lambda_{1}=0.45\pm 0.01\,\textup{km.w.e.}$ and $\lambda_{2}=0.87\pm 0.02\,\textup{km.w.e.}$, which have been determined using experimental data [mei_muon-induced_2006]. The slant depth $h$ at SNO+ is given by $h=6.01\,\textup{km.w.e.}/\cos{\theta}$. The energy of these muons is described by
@@
where $A$ is the normalisation constant, $b=0.4\,/\textup{km.w.e.}$, $\gamma_{\mu}=3.77$ and $\epsilon_{\mu}=693\,\textup{GeV}$ [mei_muon-induced_2006, groom_muon_2001].
@@ IV Event Reconstruction
The PMTs that are initially hit, as the muon enters the detector, provide information on the entry point. The muon’s entry position is found by projecting the average position of the early hit PMTs onto the inner PSUP. As the muon exits the detector, the cone of Cherenkov light is focussed around the muon’s exit point. The muon’s exit position is estimated using the charge-weighted average of all hit PMTs. The direction of the muon is taken as the vector between these two points. More details on this procedure can be found in [liggins_cosmic_2020].
@@
While the above describes the method for reconstructing muons, the reconstruction of the neutron capture signal follows a different method. The $2.2\,\textup{MeV}$ gamma ray from the neutron capture will Compton-scatter to produce secondary electrons, which produce Cherenkov radiation. The electrons are considered to be point-like events because of their relatively short track lengths, compared to the detector resolution. A likelihood approach is used to obtain the most probable position of the event, based on the PMT hit times, utilising the spherical geometry of the detector.
@@ V Data Selection
This analysis uses data taken during the water phase of SNO+, both before and after the installation of the cover gas system. The total run time across this period is $321\,\textup{days}$. Muon and neutron events are selected from the data using a suite of specifically designed selection criteria. Many of the selection criteria have been optimised using simulation with the goal of prioritising purity over efficiency, as presented in [liggins_cosmic_2020].
@@
All cosmogenic muons will travel through the cavity water, producing Cherenkov light that can be detected by the OWLs. (a) All selected cosmogenic muons must have $5$ or more OWL hits. (b) Muons produce a large amount of Cherenkov light, so events with at least $1238$ calibrated PMT hits are selected. This cut was based on the spectrum measured by SNO for cosmogenic muons, after high level cuts had been applied [collaboration_measurement_2009].
@@
The neutron selection criteria are summarised in Table 2. The neutron capture signal used for identifying neutrons creates a relatively low number of PMT hits. (a) Neutron follower events are selected if they satisfy $10\,\mu\textup{s}\leq\Delta T_{\mu}\leq 776\,\mu\textup{s}$, where $\Delta T_{\mu}$ is the time since muon. (b) Of these events, those with in-time PMT hits less than $9$ and more than $25$ are removed. The lower timing and in-time hit conditions are to account for detector ringing, an effect that is often associated with a muon, when a large charge deposit results in a temporary effective shift in the trigger threshold. Further criteria are applied to the positions of the neutrons to reduce background contributions. (c) Only neutron captures that have a perpendicular distance from their associated muon track of less than $4.6\,\textup{m}$ are selected, in order to reduce the contamination from random backgrounds within the detector. (d) Finally, the neutron capture position must have a radius of less than $8\,\textup{m}$, where the radius is the distance from the centre of the PSUP, which reduces the amount of external background contamination.
@@
where $b$ is the flat background component, $A$ is the amplitude, and $t_{n}$ is the neutron capture time, which has been set to the neutron capture time measured during the Am-Be water calibration [collaboration_measurement_2020].
@@
Figure 5: The time between the muon and neutron capture event (black), fitted with the neutron capture time measured during the Am-Be calibration (red) [collaboration_measurement_2020].
@@ VI.3 Backgrounds
One potential contribution to the background comes from other spallation products produced by the muon which mimic the neutron capture signal. This was evaluated using spallation isotope yields deduced from simulation [li_first_2014], which are compared with data in [collaboration_first_2016]. The probabilities of each isotope passing the neutron selection criteria of this analysis were applied to these yields in order to find the expected number of events per muon. Across all cosmogenic isotopes, the total number of expected background events per muon is $\mathcal{O}(10^{-7}$), which is considered to be negligible [liggins_cosmic_2020].
The other contribution comes from random background events that occur in coincidence with a muon. The rate of events which pass the neutron selection criteria is evaluated over approximately hour-long periods, outside each cosmogenic muon’s time window. This rate is then multiplied by the neutron selection time window of $766\,\mu\textup{s}$ to give the number of events expected during the selected event window following each muon. The total number of random background events, across all muons, was found to be $150\pm 2(\textup{stat})$.
@@ VI.4 Neutron Selection Efficiency
The neutron selection efficiency is defined as the fraction of neutrons produced along a section of the muon track within the detector that are detected and pass the neutron selection criteria. Simulation can be used to evaluate this fraction. For each neutron generated in simulation, the probability of selecting that neutron can be calculated, after taking into account reconstruction uncertainties and effects due to the detector response.
Am-Be calibration data has demonstrated that the efficiency for detecting neutrons depends on position within the detector [collaboration_measurement_2020]. For events close to the centre of the detector, light from the neutron capture signal has to travel further and is more susceptible to scattering. Scattering leads to a larger position reconstruction uncertainty, with some events even failing reconstruction. In this region, the efficiency is expected to be lower, and to gradually rise moving away from the centre. This slow rise ends, and the efficiency starts to decrease near the PMTs due to the lower angular acceptance of the PMTs at large incidence angles. The lower angular acceptance of the PMTs leads to a very small number of neutrons being reconstructed at large radii. Moreover, effects due to the fiducial volume start to come into play at large radii. Since each muon track is unique, spanning different lengths and regions of the detector, the neutron selection efficiency must be calculated for each muon in data.
Figure 7 shows the neutron selection efficiencies binned according to the reconstructed muon track length, calculated using $10\,\%$ of the muons selected from the data. The value of each bin is found using the weighted average of all efficiencies within the bin. The uncertainty of each bin is given by the spread of the efficiencies within the bin. Each muon in the data is assigned a neutron selection efficiency and uncertainty based on the bin corresponding to its reconstructed track length [dixon_cosmic_2025].
@@
To investigate the possible impact of the observed data-MC differences in Figures 8 and 9 on the overall neutron selection efficiency, the value of those parameters were scaled event-by-event in the MC, the neutron selection criteria re-applied, and the distributions re-generated. This allowed scaling factors that optimize the data-MC agreement to be determined, and the change in the MC neutron selection efficiency between the scaled and the un-scaled cases to be evaluated. For the perpendicular distance traversed by neutrons, a scale factor of $1.07\pm 0.03$ is preferred. Applying this scale factor results in a neutron selection efficiency that is reduced by $4-11\,\%$. This scale factor is referred to as the Distance Correction. For the number of in-time PMT hits, a scale factor of $0.95\pm 0.01$ is preferred by the muon-induced neutrons, resulting in a reduction of approximately $14\,\%$ in the neutron selection efficiency. This scale factor is referred to as the Cosmogenic Hits correction. In addition, this method is applied to the number of in-time PMT hits produced by neutron captures from the Am-Be calibration data. Using the Am-Be data significantly increases the number of events. The Am-Be data prefers a scale factor of $0.99\pm 0.02$, which causes an approximate $3\,\%$ reduction in the neutron selection efficiency. This scale factor is referred to as the Am-Be Hits correction. Table 3 lists the binned neutron selection efficiencies, from Figure 7, and the resulting neutron selection efficiencies after applying the above corrections. The use of the corrections to the neutron selection efficiency in the final yield calculation will be discussed in Section VII.
Table 3: The neutron selection efficiencies for each muon tack length bin, before and after applying the various corrections, as described in the text.

### v3

@@ I Introduction
Cosmogenic neutron production arises from a variety of processes. The muons themselves can produce neutrons through direct interactions with nuclei, known as spallation. As muons propagate through matter, they lose energy via ionisation and radiative processes, producing secondary particles that induce electromagnetic and hadronic showers. Neutrons are generated both in photonuclear interactions within the electromagnetic showers, and through spallation processes involving secondary hadrons such as protons, pions, and other nuclear fragments [1, 2].
The number of neutrons associated with a muon depends on the muon’s energy. Many underground experiments, with varying average muon energies, have measured the cosmogenic neutron yield. The relation between muon energy and neutron production was most recently demonstrated by Daya Bay [3], using liquid scintillator-based measurements [4, 5, 6, 7, 8, 9].
The average muon energy in an underground laboratory is dictated by the overburden. To date, the neutron yield measurement with the highest average muon energy was made by the Liquid Scintillator Detector (LSD) experiment [10]. Measurements using muons of comparable average energies have since been made by the Sudbury Neutrino Observatory (SNO) experiment [11] and at the China Jinping Underground Laboratory [12]. The large overburden at SNO+ means the measurement in this paper will contribute to the high energy muon data.
The medium through which a muon propagates could also play a role in the number of neutrons produced. Many measurements of the neutron yield have been performed with liquid scintillator [3, 10, 12, 7, 5, 4, 6, 9, 8]. SNO provided consistent results using both pure heavy water ($\textup{D}_{2}\textup{O}$) and NaCl-loaded heavy water. The first measurement using $\textup{H}_{2}\textup{O}$ came from Super-Kamiokande (SK), which used gadolinium-loaded water. The measurements by SK [13], in water, and KamLAND [7], in scintillator, were made using the same muon flux, providing a unique comparison of neutron production in different media. Their results were found to be consistent when normalised by medium density [13].
@@ II Experimental Set-up
SNO+ is a neutrino experiment that will search for neutrinoless double beta decay. Other physics goals of SNO+ include measurements of reactor, geo- and solar neutrinos. The detector infrastructure was inherited from the SNO experiment. While the SNO+ detector is currently filled with liquid scintillator, it was initially filled with ultra-pure water. During this period, known as the water phase, the electronics from SNO were re-commissioned and the background rates evaluated. The water phase was used as an opportunity to run the detector as a low-background water Cherenkov experiment. The water phase lasted from May 2017 to July 2019. Analyses from this phase include measurements of the $^{8}\textup{B}$ solar neutrino flux [14, 15], nucleon decay searches [16, 17], and the first evidence of detecting reactor antineutrinos in a water Cherenkov detector [18].
@@ II.1 The SNO+ Detector
The SNO+ detector, shown in Figure 1, is located $2\,\textup{km}$ ($6010\,\textup{m.w.e.}$ in water equivalent meters) underground at SNOLAB [19, 20]. An acrylic vessel (AV), of radius $6\,\textup{m}$ and thickness $5.5\,\textup{cm}$, contains the target medium that defines the phase. During the water phase, the AV was filled with $905\,\textup{tonnes}$ of ultra-pure water (UPW). A $7\,\textup{m}$ tall cylindrical neck is connected to the top of the AV, through which the AV can be filled and calibration sources can be deployed. At the top and facing down the neck are $4$ neck photomultiplier tubes (PMTs).
@@
Figure 1: Diagram of the SNO+ detector with labelled components [21].
Part way through the water phase in September 2018, a cover gas system was installed to reduce the rate of $^{222}\textup{Rn}$ induced backgrounds in the AV. For this analysis, these decays contribute to the accidental coincidence backgrounds. This system uses nitrogen gas, which acts as a barrier to radon in the air above the liquid in the neck. A detailed description of the SNO+ detector and its cover gas systems can be found in [21].
@@ II.2 Event Triggers
Each PMT has an adjustable channel threshold. When this threshold is exceeded, a hit is registered for that PMT, and a fixed-amplitude pulse of width $89\,\textup{ns}$ is produced. These pulses are summed across all PMTs. An event is recorded if the sum of the PMT pulses surpasses the event trigger threshold. For the water phase, the trigger threshold was set to $7$ pulses, approximately equivalent to a $1.4\,\textup{MeV}$ electron at the centre of the detector [22]. The trigger system is modelled in SNO+ simulations, as described in Section III.
The trigger efficiency is the probability that an event will trigger the detector, given a number of PMT hits. The time distribution of the summed PMT pulses can depend on the nature of the event. If the PMTs involved in an event all register hits at similar times, the sum of the pulses will form a sharper, narrower peak than if the PMT hits were more spread over time. Close to the threshold, this event-dependence can affect whether an event triggers the detector or not. To minimise the impact of the trigger’s dependency on the PMT hit time distribution, the trigger efficiency is calculated as a function of ‘in-time hits’, the number of PMT hits that arrive within an $89\,\textup{ns}$ time window, centred on the peak of the summed PMT pulses. In [22], it was found using simulation that the trigger efficiency reaches $100\,\%$ at or above 8 PMT hits, with reduced efficiency below that threshold. For a 2.2 MeV gamma ray, the total efficiency was found to be approximately $50\,\%$ [22].
@@ II.3 Am-Be Calibration
The results of this calibration were presented in [22]. The efficiency for detecting neutrons was found to be approximately $50\,\%$ near the centre of the detector. The neutron capture time was also measured to be $202.53^{+0.87}_{-0.76}\,\mu\textup{s}$. This neutron detection efficiency is relatively high for a water Cherenkov detector, allowing for the unique measurement of cosmogenic neutrons in UPW.
@@ III Simulation
For this analysis, Monte Carlo (MC) simulations were performed for characterising the reconstruction of muons and neutrons, and measuring the efficiency for detecting and selecting neutrons. Simulations in SNO+ are handled by RAT, a software package used for offline analysis [23]. All low level physics simulation and particle tracking is performed using the GEANT4 (version 10.00.p04) software package [24], which has been incorporated into RAT, along with the detector geometry and trigger performance. Calibrations have been performed in order to tune and verify the full SNO+ simulation model.
@@
where $\theta$ is the angle from zenith from which the muon arrives. The parameters used are $I_{1}=(8.60\pm 0.53)\times 10^{-6}\,\textup{cm}^{-2}\textup{s}^{-1}$, $I_{2}=(0.44\pm 0.06)\times 10^{-6}\,\textup{cm}^{-2}\textup{s}^{-1}$, $\lambda_{1}=0.45\pm 0.01\,\textup{km.w.e.}$ and $\lambda_{2}=0.87\pm 0.02\,\textup{km.w.e.}$, which have been determined using experimental data [25]. The slant depth $h$ at SNO+ is given by $h=6.01\,\textup{km.w.e.}/\cos{\theta}$. The energy of these muons is described by
@@
where $A$ is the normalisation constant, $b=0.4\,/\textup{km.w.e.}$, $\gamma_{\mu}=3.77$ and $\epsilon_{\mu}=693\,\textup{GeV}$ [25, 26].
@@ IV Event Reconstruction
The inward-looking PMTs that are initially hit, as the muon enters the detector, provide information on the entry point. The muon’s entry position is found by projecting the average position of the early hit PMTs onto the inner PSUP. As the muon exits the detector, the cone of Cherenkov light is focussed around the muon’s exit point. The muon’s exit position is estimated using the charge-weighted average of all hit PMTs. The direction of the muon is taken as the vector between these two points. More details on this procedure can be found in [27].
@@
While the above describes the method for reconstructing muons, the reconstruction of the neutron capture signal follows a different method. The $2.2\,\textup{MeV}$ gamma ray from the neutron capture will Compton-scatter to produce secondary electrons, which produce Cherenkov radiation. These electrons have up to $1\,\textup{cm}$ track lengths in water, which is small compared to the $\sim 20\,\textup{cm}$ position reconstruction resolution for low-energy events in the SNO+ water phase [16]. The neutron capture signal is therefore treated as a point-like event. A likelihood approach is used to obtain the most probable event position, based on PMT hit times, utilising the spherical geometry of the detector. For an assumed event vertex, the distribution of PMT hit time residuals (the difference between the measured PMT hit times and the expected photon arrival times) is calculated. The likelihood is constructed by comparing the observed distribution of time residuals to the expected distribution, obtained from simulation, and is maximised with respect to the event position.
@@ V Data Selection
This analysis uses data taken during the water phase of SNO+, both before and after the installation of the cover gas system. The total run time across this period is $321\,\textup{days}$. Muon and neutron events are selected from the data using a suite of specifically designed selection criteria. Many of the selection criteria have been optimised using simulation with the goal of prioritising purity over efficiency, as presented in [27].
@@
All cosmogenic muons will travel through the cavity water, producing Cherenkov light that can be detected by the OWLs. (a) All selected cosmogenic muons must have $5$ or more OWL hits. (b) Muons produce a large amount of Cherenkov light, so events with at least $1238$ calibrated PMT hits are selected. This cut was based on the spectrum measured by SNO for cosmogenic muons, after high level cuts had been applied [28].
@@
The neutron selection criteria are summarised in Table 2. The neutron capture signal used for identifying neutrons creates a relatively low number of PMT hits. (a) Neutron follower events are selected if they satisfy $10\,\mu\textup{s}\leq\Delta T_{\mu}\leq 776\,\mu\textup{s}$, where $\Delta T_{\mu}$ is the time since muon. (b) Of these events, those with in-time PMT hits less than $9$ and more than $25$ are removed. The lower timing and in-time hit conditions are to account for detector ringing, an effect that is often associated with a muon, when a large charge deposit results in a temporary effective shift in the trigger threshold. Further criteria are applied to the positions of the neutrons to reduce background contributions, which have been optimised through simulation. (c) Only neutron captures that have a perpendicular distance from their associated muon track of less than $4.6\,\textup{m}$ are selected, in order to reduce the contamination from random backgrounds within the detector. (d) Finally, the neutron capture position must have a radius of less than $8\,\textup{m}$, where the radius is the distance from the centre of the PSUP, which reduces the amount of external background contamination.
@@
where $b$ is the flat background component, $A$ is the amplitude, and $t_{n}$ is the neutron capture time, which has been set to the neutron capture time measured during the Am-Be water calibration [22].
@@
Figure 5: The time between the muon and neutron capture event (black), fitted with the neutron capture time measured during the Am-Be calibration (red) [22].
@@ VI.3 Backgrounds
One potential contribution to the background comes from other spallation products produced by the muon which mimic the neutron capture signal. This was evaluated using spallation isotope yields deduced from simulation [29], which are compared with data in [30]. The probabilities of each isotope passing the neutron selection criteria of this analysis were applied to these yields in order to find the expected number of events per muon. Across all cosmogenic isotopes, the total number of expected background events per muon is $\mathcal{O}(10^{-7}$), which is considered to be negligible [27].
The other contribution comes from random background events, from internal and external radioactivity, that occur in coincidence with a muon. The rate of events which pass the neutron selection criteria is evaluated over approximately hour-long periods, outside each cosmogenic muon’s time window. This rate is then multiplied by the neutron selection time window of $766\,\mu\textup{s}$ to give the number of events expected during the selected event window following each muon. The total number of random background events, across all muons, was found to be $150\pm 2(\textup{stat})$.
@@ VI.4 Neutron Selection Efficiency
The neutron selection efficiency is defined as the fraction of neutrons produced along a section of the muon track within the detector that are detected and pass the neutron selection criteria. Simulation can be used to evaluate this fraction. For each neutron generated in simulation, the probability of selecting that neutron can be calculated, after taking into account reconstruction uncertainties and effects due to the detector response, including trigger efficiency.
Am-Be calibration data has demonstrated that the efficiency for detecting neutrons depends on position within the detector [22]. For events close to the centre of the detector, light from the neutron capture signal has to travel further and is more susceptible to scattering. Scattering leads to a larger position reconstruction uncertainty, with some events even failing reconstruction. In this region, the efficiency is expected to be lower, and to gradually rise moving away from the centre. This slow rise ends, and the efficiency starts to decrease near the PMTs due to the lower angular acceptance of the PMTs at large incidence angles. The lower angular acceptance of the PMTs leads to a very small number of neutrons being reconstructed at large radii. Moreover, effects due to the fiducial volume start to come into play at large radii. Since each muon track is unique, spanning different lengths and regions of the detector, the neutron selection efficiency must be calculated for each muon in data.
Figure 7 shows the neutron selection efficiencies binned according to the reconstructed muon track length, as defined in Section VI.2, calculated using $10\,\%$ of the muons selected from the data. The value of each bin is found using the average of all efficiencies within the bin, weighted by their statistical uncertainties. The uncertainty of each bin is given by the spread of the efficiencies within the bin. Each muon in the data is assigned a neutron selection efficiency and uncertainty based on the bin corresponding to its reconstructed track length [31].
@@
To investigate the possible impact of the observed data-MC differences in Figures 8 and 9 on the overall neutron selection efficiency, the value of those parameters were scaled event-by-event in the MC, the neutron selection criteria re-applied, and the distributions re-generated. This allowed scaling factors that optimize the data-MC agreement to be determined, and the change in the MC neutron selection efficiency between the scaled and the un-scaled cases to be evaluated. For the perpendicular distance traversed by neutrons, a scale factor of $1.07\pm 0.03$ is preferred. Applying this scale factor results in a neutron selection efficiency that is reduced by $4-11\,\%$. This scale factor is referred to as the Distance Correction. For the number of in-time PMT hits, a scale factor of $0.95\pm 0.01$ is preferred by the muon-induced neutrons, resulting in a reduction of approximately $14\,\%$ in the neutron selection efficiency. This scale factor is referred to as the Cosmogenic Hits correction. In addition, this method is applied to the number of in-time PMT hits produced by neutron captures from the Am-Be calibration data. Using the Am-Be data significantly increases the number of events. The Am-Be data prefers a scale factor of $0.99\pm 0.02$, which causes an approximate $3\,\%$ reduction in the neutron selection efficiency. This scale factor is referred to as the Am-Be Hits correction. Table 3 lists the binned neutron selection efficiencies before corrections, from Figure 7, and the resulting neutron selection efficiencies after applying the above corrections. The use of the corrections to the neutron selection efficiency in the final yield calculation will be discussed in Section VII.
Table 3: The neutron selection efficiencies for each muon track length bin, before and after applying the various corrections, as described in the text.